Shift summary for evening 7/17/99

Jason Hamm (jhamm@physics.arizona.edu)
Sat, 17 Jul 1999 22:18:32 -0700 (MST)


Shift Summary for evening shift, 07/17/99

On Shift: J. Whitmore, J. Hamm

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.25
Hours of E832 data (special runs not included): 5.75
Hours of special runs (laser scans, muon runs...): 5 min

Accounting of lost time: DA problems
========================================================================
Runs ended this shift:

Run 13934 - Beam 832 (14.8M events)
Run 13936 - 4 E12 Muon (~100K events)

Tapes brought to FCC this shift: QK*146-148, QKB050
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: 832 data taking at ~4.2 E12,with a break for a muon run
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13934 88 16:45 HV mainframe 5 turns off for one spill while
the CSI HV's are being verified via hv1440.exe.
Reason still unknown - no problems since.

13934 150 17:00 Crate 53 busy - fixed by kfrend

18:00 Dave N. installs 3 new bottles of argon-ethane
on the west side of the gas hut. Helium tubes
will be changed over tomorrow morning.

13934 170 18:32 Missing FFFF word in stream 5 - fixed by kfrend

13934 213 19:30 SEM drops to ~2 E12 and BOA rises from 1.2 to 1.4.
MCR says they're working on tuning. Back to
normal at spill 223.

21:09 Tapes fill up and run stops. Prepare for muon run.

21:33 Start muon run 13936

21:36 Stop muon run and ask MCR to set us up for hadrons
again. NM2-3S comes up to only 320A, though
(nominal is 875A). MCR claims this what we're
SUPPOSED to have for data mode...eventually, they
are persuaded to bring it up to 875A and we
initialize Beam 832 run 13937. Rick F. and Rick
C. are notified.

21:50 Start run 13937

13937 5 21:55 DDD4 veto enabled. Not cleared by updating busy
logic. Stop run, but stop process hangs. Try to
fresh start and init twice, but init hangs in
stage 3. Problem looks to be in stream 5 of plane
4 - no heartbeat. Pushed reset button & try again
- still hangs.

Paged Jason L, who advises us to power cycle the
memory module. We do so and heartbeat for stream
5 returns.

23:14 Run 13941 successfully initialized and started!!!
Since there was only ~5 spills of data on the
tapes before the DA problem, we just overwrote.
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

AZLAT in crate 53, slot 13 should be replaced - page Jason H. during next
significant non-owl downtime.